# Investigating Emotions, Social Roles, Well-being and Mobile Usage Behaviors During COVID-19 Home Confinement

### v1

@@ 6.2. Prediction Pipeline
while emotion prediction can be addresed as a classification problem, with emotion levels categorized into two or three classes based on predefined thresholds (65; 93). We believe that utilizing regression is a better approach becuase valence and arousal labels are inherently ordinal data, where the relative order holds significance. In this research, we establish a regression-based pipeline to forecast the emotion scores of participants, as detailed below.
Emotion. The ESM questionnaires, completed by the participants every 90 minutes, serve as the ground truth in the regression models developed in this study. Specifically, the item ‘valence’ was rated on a 5-point Likert scale from ‘unhappy’ to ‘happy’, where 1 indicates ‘unhappy’ and 5 refers to ‘happy’. Similarly, another item ‘arousal’ was also rated with a 5-point Likert scale from 1-5, with 1 indicates ‘calm’ and 5 means ‘excited’. To predict the levels of valence and arousal, the app activities within 90 minutes prior to completing the ESM questionnaire were leveraged as features.
Regressors. In this research, we aim to develop regression models that accurately predicts human emotions during home confinement. To accomplish this, we employ three commonly used regression models: K-Nearest Neighbor (KNN) (79) regressor , Random Forest (RF) (78) regressor, and Gradient Boosting (GB) (27) regressor. KNN is a widely utilized model known for its versatility across various domains. It works by finding the K nearest neighbors to a given data point and using their values to predict the target variable. RF and GB, on the other hand, are ensemble models based on decision trees. These models combine predictions from multiple trees. However, there is a distinction in their approach. GB constructs trees or aggregates results in a way thatminimizes the loss function, whereas RF constructs individual trees or aggregates results simultaneously.
@@
Validation. To compensate for the limited richness of our data, we have implemented nested cross-validation (CV) (57) to train and evaluate our prediction model. This approach allows us to accurately estimate the unbiased generalization performance of the model. The nested CV consists of an inner loop and an outer loop. Initially, the dataset is divided into training, validation, and testing sets. This helps optimize the model’s performance. In the outer loop, similar to previous human-centered studies (30; 24), we divide the data into multiple groups, each group representing one participant. To ensure unbiased evaluation, we apply k-fold cross validation (88) where k=5 on all groups. This means that the data from the same participants are not used in both the training and test sets simultaneously. In the inner loop, the remaining data groups are further split into three folds, with each fold serving as a validation set. We optimize the hyperparameters on the training set, evaluate them on the validation set, and select the best parameter settings based on the results from the three folds.
@@
(d) MAE for arousal
Figure 8. Prediction performance for valance and arousal with different feature sets, where the error bar presents a 95% confidence interval
Baselines. Similar to previous human-centered studies (30; 87), we compare the proposed model with two baselines. The first baseline is the Mean baseline, which calculates the average score of user responses. The second baseline is the Median baseline, which calculates the median value based on the data distribution and treats it as a predicted value. To evaluate the performance of regression models, we employ two commonly used metrics: Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) (18). Generally, lower values of MAE or RMSE indicate better prediction performance.
@@ 6.3.1. General Model
Table 5. Prediction performance of the general models with different feature sets
Prediction Result (RMSE/MAE)
@@
Valence (13) 0.986/0.824 1.091/0.767 0.936/0.730 0.912/0.725 0.987/0.724
Arousal (12) 1.113/0.884 1.146/0.784 1.064/0.819 1.011/0.774 1.076/0.745
Table 5 displays the prediction results of the general models for valence and arousal. Regarding the valence dimension, our results indicate that the RF regressor achieved the best predictive performance across three different feature sets. For the regression model using all features, the RMSE obtained by the RF algorithm was 0.903, surpassing the Median baseline by 17.23%. Furthermore, the MAE of the RF regressor using all feature sets was 0.717, outperforming the mean baseline by 12.99% and the median baseline by 6.52%. Similar results were observed for the models utilizing the top 17 features. Regarding the arousal dimension, our results indicate that the RF regressor utilizing the all features achieved the best performance in terms of RMSE, with a value of 1.005. This outperformed the Median baseline by 12.30%. These outcomes suggest that the selected features serve as effective predictors of arousal for participants during home confinement.
@@ 6.3.2. Individual Models
Additionally, it is worth noting that while the proposed valence models exhibited better prediction performance than baseline models for certain participants (e.g., ‘M02MMP’, ‘L10FGB’), there were participants who did not achieve satisfactory results. The possible reasons may be two-fold: firstly, individual differences play a crucial role in determining valence, and it may not be straightforward to accurately predict valence solely based on mobile usage behaviors for some participants. This highlights the complexity and subjectivity of valence and the need for a more nuanced understanding of individual emotional experiences; secondly, insufficient data availability may have limited the performance of the predictive models for some participants. Having a larger and more diverse dataset would enable better training and improve the performance of the predictive models. These findings contribute to our understanding of personalized prediction models for well-being, which could ultimately lead to more tailored interventions and support systems to enhance individuals’ well-being during periods of home confinement.
@@ 6.4. Feature Importance
Both models consistently identify ‘sin_time’ and ‘cos_time’ as the top two important features, indicating that the time period of the day significantly influences participants’ valence and arousal. Additionally, the day of the week (‘weekday’) impacts both models, although it has less significance in the arousal model. Surprisingly, the feature ‘is_weekend’ does not appear among the top 15 features, suggesting that while there is still variation in terms of time and day, weekdays may not have a significant impact on participants’ valence and arousal levels. These findings differ slightly from the observations reported in (40), which found that app-usage patterns varied greatly vary by time and day. The discrepancy may be attributed to the unique circumstances of home confinement, where individuals may experience a loss of time perception and a blurring of the distinction between weekdays and weekends. This aligns with the findings presented in (17), which suggest that the perception of time and daily routines can be disrupted during periods of confinement.
Above findings have important implications for understanding the factors that contribute to human well-being, particularly in the context of home confinement. They provide valuable insights that can inform the development of interventions aimed at promoting emotional wellness during such periods. By considering the impact of app usage, notifications, and time-related factors, we can also design targeted strategies to enhance individuals’ emotional well-being and overall quality of life.
@@ 7. Implications
Our study examines the impact of home confinement during the COVID-19 lockdown on individuals’ well-being, social roles, and work-life balance.Through the use of unobtrusive mobile phone sensing, we aim to predict participants’ well-being and social roles. The findings from our study are significant in raising awareness about the importance of mental health not only during periods of home confinement but also in similar situations such as post-operative recovery, parental leave, aging, or limited mobility related to disability. These findings can also play a crucial role in the early detection and treatment of mental illnesses that may arise during these circumstances. Additionally, policymakers can utilize our research to develop appropriate strategies that minimize the negative effects of pandemic restrictive policies on mental health. To achieve this, quarantine or isolation time should be kept as short as possible, and voluntary quarantine could be promoted to reduce long-term mental health complications, as recommended by Brooks et al. (11).
@@
In the post-pandemic era, the traditional office-only working arrangement has shifted to more flexible arrangements, with an increasing number of people opting to work from home (49). Studies have shown that individuals who continued to work from home after the pandemic exhibited similar behaviors to those during the pandemic (53), underscoring the importance of studying people’s behaviors in high technology-reliance contexts. Our study was conducted in a city with the longest home confinement period, offering optimal experimental conditions to investigate the influence of app usage on well-being while minimizing confounding effects from other variables, such as nature and physical social interactions. Furthermore, our study contributes to understanding the mental health benefits and consequences of working from home or relying heavily on technology. For instance, tracking mental well-being status provides a means of measuring work-life balance and enables individuals to manage their working schedules or arrangements (e.g., work-from-home or half-office half-home) more effectively.
@@
In addition, this study provides a valuable opportunity for researchers to study the psychological states and smartphone usage behaviors of individuals under home confinement conditions, which has important implications for various fields. For example, it may inform the development of interventions to support individuals in similar situations, such as those willing or undergoing post-operative home rehabilitation. Understanding how individuals cope with home confinement and the role that technology plays in their lives can help healthcare professionals or researchers develop effective interventions and design effective remote support programs to improve the overall quality of care for patients in similar situations. Furthermore, the findings from this study may have broader implications for our understanding of human behavior and mental health during times of crisis, informing the development of new policies and interventions to support individuals in future crises.
@@ 8. Limitations
We have identified three main limitations in this research. One significant limitation is the selection of time and place. Regarding time, we were unable to compare user well-being and digital device usage data before COVID-19 lockdown with the data during the lockdown due to the unpredictable nature of lockdown policies. However, we have mitigated this limitation by allowing users to report the impact of the lockdown on their well-being and work-life balance through self-reported questionnaires. These data clearly demonstrate the effects of lockdowns on people’s lives, as discussed in Section 4. In terms of geography, our data collection was limited to Melbourne, where the COVID-19 lockdown with Stage 4 restrictions was applied. While the specific measures taken by different countries and regions vary, and the overall mental well-being of populations may differ, studies conducted in other countries have also demonstrated the impact of COVID-19 on human well-being and mental health (10; 14; 11). Therefore, our research still holds value in informing government officials or managers to make appropriate decisions and highlighting the importance of mental health care during home confinement situations.
The second limitation is the relatively short duration of the data collection period. We chose to collect data for only three weeks in order to minimize interruptions from ESM questionnaires and end-of-day surveys, with the goal of reducing the burden on participants during the challenging realities of COVID-19 home confinement. Despite the short duration, we were still able to collect a substantial amount of data, including 502,485 records of users’ context and phone usage, 1,749 ESM responses, and 265 end-of-day survey responses. It is worth noting that the delivery of end-of-day survey link through emails allowed users to complete the survey voluntarily at their convenience. Although the number of end-of-day survey responses is small, the optional nature of the surveys, especially text questions, ensures the credibility and reliability of the collected data. However, due to the limited three-week period of our data collection, it is important to acknowledge that if strict home confinement lasts longer, the emotional experiences of individuals may evolve as they adapt to new challenges.
@@ 9. Conclusion
In this study, we delved into how human emotions, social roles, and mobile usage behaviors were affected during the COVID-19 home confinement period. Our proposed models yielded promising results in predicting human emotions during home confinement. Moreover, we uncovered intriguing insights regarding the correlation between different factors and human well-being. To address our research objectives on home confinement, we formulated three research questions and key findings are summarized below.
@@
Despite the challenges we faced in collecting data during COVID-19 lockdowns, we believe that gaining insights into human well-being and mobile usage behaviors during home confinement is crucial. This understanding is not only valuable for policymakers in implementing effective measures during future outbreaks, but also for individuals to become more self-aware of their well-being and mental health. Moving forward, our research will expand to investigate users’ desktop usage behavior. By comparing the differences in usage behavior across different platforms, we aim to gain a deeper understanding of people’s daily lives and work-life balance.
@@ 10. Acknowledgments
We sincerely thank the anonymous reviewers.

### v2

@@ 6.2. Prediction Pipeline
while emotion prediction can be addressed as a classification problem, with emotion levels categorized into two or three classes based on predefined thresholds (70; 98). We believe that utilizing regression is a better approach because valence and arousal labels are inherently ordinal data, where the relative order holds significance. In this research, we establish a regression-based pipeline to forecast the emotion scores of participants, as detailed below.
Emotion. The ESM questionnaires, completed by the participants every 90 minutes, serve as the ground truth in the regression models developed in this study. Specifically, the item ‘valence’ was rated on a 5-point Likert scale from ‘unhappy’ to ‘happy’, where 1 indicates ‘unhappy’ and 5 refers to ‘happy’. Similarly, another item ‘arousal’ was also rated with a 5-point Likert scale from 1-5, with 1 indicating ‘calm’ and 5 means ‘excited’. To predict the levels of valence and arousal, the app activities within 90 minutes prior to completing the ESM questionnaire were leveraged as features.
Regressors. In this research, we aim to develop regression models that accurately predicts human emotions during home confinement. To accomplish this, we employ three commonly used regression models: K-Nearest Neighbor (KNN) (84) regressor, Random Forest (RF) (83) regressor, and Gradient Boosting (GB) (29) regressor. KNN is a widely utilized model known for its versatility across various domains. It works by finding the K nearest neighbours to a given data point and using their values to predict the target variable. RF and GB, on the other hand, are ensemble models based on decision trees. These models combine predictions from multiple trees. However, there is a distinction in their approach. GB constructs trees or aggregates results in a way that minimizes the loss function, whereas RF constructs individual trees or aggregates results simultaneously.
@@
Validation. To compensate for the limited richness of our data, we have implemented nested cross-validation (CV) (61) to train and evaluate our prediction model. This approach allows us to accurately estimate the unbiased generalization performance of the model. The nested CV consists of an inner loop and an outer loop. Initially, the dataset is divided into training, validation, and testing sets. This helps optimize the model’s performance. In the outer loop, similar to previous human-centred studies (33; 25), we divide the data into multiple groups, each group representing one participant. To ensure unbiased evaluation, we apply k-fold cross validation (93) where k=5 on all groups. This means that the data from the same participants are not used in both the training and test sets simultaneously. In the inner loop, the remaining data groups are further split into three folds, with each fold serving as a validation set. We optimize the hyperparameters on the training set, evaluate them on the validation set, and select the best parameter settings based on the results from the three folds.
@@
(d) RMSE for arousal
Figure 8. Prediction performance for valance and arousal with different feature sets, where the error bar presents standard deviation.
Baselines. Similar to previous human-centered studies (33; 92), we compare the proposed model with two baselines. The first baseline is the Mean baseline, which calculates the average score of user responses. The second baseline is the Median baseline, which calculates the median value based on the data distribution and treats it as a predicted value. To evaluate the performance of regression models, we employ two commonly used metrics: Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) (19). Generally, lower values of MAE or RMSE indicate better prediction performance.
@@ 6.3.1. General Model
Table 5. Prediction performance of the general models for valence and arousal with different feature sets. The regression models are K-Nearest Neighbor (KNN), Random Forest (RF) and Gradient Boosting (GB), compared with Mean and Median baseline. The evaluation metrics are RMSE and MAE (with standard deviation).
Prediction Result (RMSE(SD)/MAE(SD))
@@
[…]
Table 5 displays the prediction results of the general models for valence and arousal. Regarding the valence dimension, our results indicate that the RF regressor achieved the best predictive performance across three different feature sets. For the regression model using all features, the RMSE obtained by the RF algorithm was 0.903, surpassing the Median baseline by 17.23%. Furthermore, the MAE of the RF regressor using all feature sets was 0.717, outperforming the mean baseline by 12.99% and the median baseline by 6.52%. Similar results were observed for the models utilizing the top 17 features. Regarding the arousal dimension, our results indicate that the RF regressor utilizing all features achieved the best performance in terms of RMSE, with a value of 1.005. This outperformed the Median baseline by 12.30%. These outcomes suggest that the selected features serve as effective predictors of arousal for participants during home confinement.
@@ 6.3.2. Individual Models
Additionally, it is worth noting that while the proposed valence models exhibited better prediction performance than baseline models for certain participants (e.g., ‘M02MMP’, ‘L10FGB’), there were participants who did not achieve satisfactory results. The possible reasons may be two-fold: firstly, individual differences play a crucial role in determining valence, and it may not be straightforward to accurately predict valence solely based on mobile usage behaviours for some participants. This highlights the complexity and subjectivity of valence and the need for a more nuanced understanding of individual emotional experiences; secondly, insufficient data availability may have limited the performance of the predictive models for some participants. Having a larger and more diverse dataset would enable better training and improve the performance of the predictive models. These findings contribute to our understanding of personalized prediction models for well-being, which could ultimately lead to more tailored interventions and support systems to enhance individuals’ well-being during periods of home confinement.
@@ 6.4. Feature Importance
Both models consistently identify ‘sin_time’ and ‘cos_time’ as the top two important features, indicating that the time period of the day significantly influences participants’ valence and arousal. Additionally, the day of the week (‘weekday’) impacts both models, although it has less significance in the arousal model. Surprisingly, the feature ‘is_weekend’ does not appear among the top 15 features, suggesting that while there is still variation in terms of time and day, weekdays may not have a significant impact on participants’ valence and arousal levels. These findings differ slightly from the observations reported in (44), which found that app usage patterns varied greatly by time and day. The discrepancy may be attributed to the unique circumstances of home confinement, where individuals may experience a loss of time perception and a blurring of the distinction between weekdays and weekends. This aligns with the findings presented in (18), which suggest that the perception of time and daily routines can be disrupted during periods of confinement.
The above findings have important implications for understanding the factors that contribute to human well-being, particularly in the context of home confinement. They provide valuable insights that can inform the development of interventions aimed at promoting emotional wellness during such periods. By considering the impact of app usage, notifications, and time-related factors, we can also design targeted strategies to enhance individuals’ emotional well-being and overall quality of life.
@@ 7. Implications
Our study examines the impact of home confinement during the COVID-19 lockdown on individuals’ well-being, social roles, and work-life balance. Through the use of unobtrusive mobile phone sensing, we aim to predict participants’ well-being and social roles. First of all, our findings highlight the importance of mental health during periods of home confinement. Home confinement is not limited to temporary pandemic adaptation but also the growing trend of working from home post-pandemic (40; 69) and other similar situations, such as post-operative recovery, parental leave, or limited mobility related to disability, or even mobile phone addiction. Secondly, our quantitative approach and regression models can contribute to tracking, early detection and treatment of mental illnesses during these circumstances. Our results also help policymakers to develop appropriate strategies that minimize the long-term negative effects of the pandemic on mental health and encourage people to build healthy digital habits (69). For example, quarantine or isolation time should be kept as short as possible, and voluntary quarantine could be promoted to reduce long-term mental health complications, as recommended by Brooks et al. (12).
@@
In the post-pandemic era, the traditional office-only working arrangement has shifted to more flexible arrangements, with an increasing number of people opting to work from home (53; 40). Studies have shown that individuals who continued to work from home after the pandemic exhibited similar behaviours to those during the pandemic (57), underscoring the importance of studying people’s behaviours in high technology-reliance contexts. Our study was conducted in a city with the longest home confinement period, offering optimal experimental conditions to investigate the influence of app usage on well-being while minimizing confounding effects from other variables, such as nature and physical social interactions. Furthermore, our study contributes to understanding the mental health benefits and consequences of working from home or relying heavily on technology. For instance, tracking mental well-being status provides a means of measuring work-life balance and enables individuals to manage their working schedules or arrangements (e.g., work-from-home or half-office half-home) more effectively.
@@
In addition, this study provides a valuable opportunity for researchers to study the psychological states and smartphone usage behaviours of individuals under home confinement conditions, which has important implications for various fields. For example, it may inform the development of interventions to support individuals in similar situations, such as those willing or undergoing post-operative home rehabilitation. Rethinking about the future of work and life scenarios, where heavily rely on digital technologies. It is important to understand how people deal with being confined to their homes and how technology impacts their lives. This understanding can help researchers design smart environments, or assist healthcare professionals in developing effective remote support programs that enhance the quality of care for those in similar situations. Furthermore, the findings from this study may have broader implications for our understanding of human behaviour and mental health during times of crisis, informing the development of new policies and interventions to support individuals in future crises.
@@ 8. Limitations
We have identified three main limitations in this research. One significant limitation is the selection of time and place. Regarding time, we were unable to compare user well-being and digital device usage data before COVID-19 lockdown with the data during the lockdown due to the unpredictable nature of lockdown policies. However, we have mitigated this limitation by allowing users to report the impact of the lockdown on their well-being and work-life balance through self-reported questionnaires. These data clearly demonstrate the effects of lockdowns on people’s lives, as discussed in Section 4. In terms of geography, our data collection was limited to Melbourne, where the COVID-19 lockdown with Stage IV restrictions was applied. While the specific measures taken by different countries and regions vary, and the overall mental well-being of populations may differ, studies conducted in other countries have also demonstrated the impact of COVID-19 on human well-being and mental health (11; 15; 12). Therefore, our research still holds value in informing government officials or managers to make appropriate decisions and highlighting the importance of mental health care during home confinement situations.
The second limitation is the relatively short duration of the data collection period. We chose to collect data for only three weeks in order to minimize interruptions from ESM questionnaires and end-of-day surveys, with the goal of reducing the burden on participants during the challenging realities of COVID-19 home confinement. Despite the short duration, we were still able to collect a substantial amount of data, including 502,485 records of users’ context and phone usage, 1,749 ESM responses, and 265 end-of-day survey responses. It is worth noting that the delivery of the end-of-day survey link through emails allowed users to complete the survey voluntarily at their convenience. Although the number of end-of-day survey responses is small, the optional nature of the surveys, especially text questions, ensures the credibility and reliability of the collected data. However, due to the limited three-week period of our data collection, it is important to acknowledge that if strict home confinement lasts longer, the emotional experiences of individuals may evolve as they adapt to new challenges.
@@ 9. Conclusion
In this study, we delved into how human emotions, social roles, and mobile usage behaviours were affected during the COVID-19 home confinement period. Our proposed models yielded promising results in predicting human emotions during home confinement. Moreover, we uncovered intriguing insights regarding the correlation between different factors and human well-being. To address our research objectives on home confinement, we formulated three research questions and key findings are summarized below.
@@
Despite the challenges we faced in collecting data during COVID-19 lockdowns, we believe that gaining insights into human well-being and mobile usage behaviours during home confinement is crucial. This understanding is not only valuable for policymakers in implementing effective measures during future outbreaks, but also for individuals to become more self-aware of their well-being and mental health. Moving forward, our research will expand to investigate users’ desktop usage behaviour. By comparing the differences in usage behaviour across different platforms, we aim to gain a deeper understanding of people’s daily lives and work-life balance.
@@ 10. Acknowledgments
We sincerely thank the anonymous reviewers. We state that this manuscript has no relation to our prior publications (or concurrently submitted papers).